### v1

Abstract
Robot navigation is increasingly crucial across applications like delivery services and warehouse management. The integration of Reinforcement Learning (RL) with classical planning has given rise to meta-planners that combine the adaptability of RL with the explainable decision-making of classical planners. However, the exploration capabilities of RL based meta-planners during training are often constrained by the capabilities of the underlying classical planners. This constraint can result in limited exploration, thereby leading to sampling skew issues. To address these issues, our paper introduces a novel framework, Digimon, which begins with behavior-guided diagnosis for exploration bottlenecks within the meta-planner and follows up with a mitigation strategy that conducts up-sampling from diagnosed bottleneck data. Our evaluation shows 13.5%+ improvement in navigation performance, greater robustness in out-of-distribution environments, and a 4$\times$ boost in training efficiency. Digimon is designed as a versatile, plug-and-play solution, allowing seamless integration into various RL-based meta-planners.
I Introduction
Robot navigation is essential for applications like delivery services, hospital logistics, warehouse management, and library automation. Navigating cluttered environments remains a key challenge for achieving agile, efficient, and safe movement. Classical planners [abdulsaheb2022robot, asano1985visibility, Canny1985AVM, Sampling2014] address this by using specific cost functions and rules, which are difficult to generalize and often require expert manual tuning [Campbell2020, raheem2022robot, zheng2021rostuning, feng2024rocas]. In contrast, learning-based approaches [pfeiffer2017perception, gao2017intention, chiang2019learning, xiao2021toward, liu2021lifelong, xiao2021agile, xiao2020appld] offer adaptability but can suffer from instability and lack interpretability, leading to erratic behaviors that compromise safety and efficiency.
To overcome these limitations, recent studies introduced meta-planners [xu2021applr, xiao2022motion] that integrate RL with classical planning frameworks (see Fig. 1). RL-based meta-planners dynamically adjust the parameters of classical planners, combining RL’s adaptability with the explainable decision-making of classical methods for balanced robot navigation.
However, existing meta-planners suffer from sampling skew due to the integration of RL and classical planners. The exploration capabilities of RL are often limited by the classical planner’s inherent strategies. For instance, if the classical planner prefers conservative paths, the RL meta-planner may be unable to explore more aggressive or unconventional routes, resulting in sub-optimal performance. Previous approaches [swazinna2021overcoming, shen2023loose, mark2024offline, Strehl2005ATA, nikishin2022primacy], such as adding noise to RL actions, do not address these specific integration challenges. Therefore, there is a crucial need for specialized methods to mitigate the biases introduced by combining RL meta-planners with classical planning.
(a) Classical
(b) RL End-to-End
(c) Hybrid
Figure 1: Different Types of Planners. Note that the meta-planner is applied onto classical local planner $f$, not global trajectory planner.
In this paper, we propose Digimon, a diagnosis and mitigation framework addressing the limitations of RL meta-planners. Digimon starts with a behavior-guided diagnostic phase to identify exploration bottlenecks caused by classical planners. It then employs a novel up-sampling strategy to sample data from these bottleneck areas, providing RL meta-planner with a balanced replay buffer. This enhances RL exploration ability and mitigates sampling skew caused by classical planner’s preference. In summary, our contributions are:
  • We identify and analyze the exploration bottlenecks and sampling skews arising from integrating RL with classical planning in meta-planners.
  • We present Digimon, featuring a diagnostic phase to detect bottlenecks and a mitigation phase that uses data up-sampling to improve RL meta-planner training.
  • Digimon achieves 13.5%+ improvement in navigation performance, greater robustness in out-of-distribution environments, and a 4$\times$ increase in training efficiency. Additionally, Digimon is a versatile, plug-and-play solution, compatible with various RL-based meta-planners.
II Background, Notations and Related Work
II-A Motion Planning
Motion planning is used to determine a trajectory for a robot to move from a start position to a goal position while avoiding obstacles (as visualized in Fig. 2). It is worth noting that there are 2 major categories of planning algorithms: local trajectory planners (e.g. DWA [fox1997dwa], TEB [teb]) and global path planners (e.g., Dijkstras [dijkstra], A* [astar], RRT [lavalle1998rrt, lavalle2001rrt]).
In this paper, we consider motion planning in navigation tasks and assume that the robot employs a classical local motion planner $f$. The local planner $f$ can be tuned via a set of configurations (i.e., configurable parameters) $\theta\in\Theta$, where $\Theta$ denotes all possible values of planner configuration. While navigating in a physical world $\mathcal{W}$ with obstacles, $f$ tries to move the robot to a local goal $g_{t}\in\mathbb{R}^{2}$ (computed by the global planner). At each time step $t$, $f$ receives sensor observations $o_{t}$ (e.g. lidar scans) to do collision avoidance, and then attempts to reach the local goal $g_{t}$ in a fast and collision-free manner. The local planner $f$ is responsible for computing the motion commands $u_{t}=f(o_{t},g_{t}|\theta)$ (e.g., angular or linear velocity) at each time step to reach $g_{t}$.
II-B Meta-Planner
Figure 2: Maps of various difficulty levels.
Meta-Planner operates on top of the original planner $f$, adjusting its configuration $\theta$. Fig. 1c shows that a hybrid planner with a meta-planner combines the explainability of classical planners with the adaptability of RL-based planners, enhancing performance. It has been applied across domains such as robot navigation [fridovich2018planning, xiao2022motion, siva2019robot, pokle2019deep], autonomous driving [jiang2022efficient], drone control [loquercio2022autotune], and robotic manipulation [Moll2021HyperPlanAF, jia2023dynamic].
There are two types of meta-planners: static and dynamic. Static meta-planners tune the configuration $\theta$ before deployment, while dynamic meta-planners adjust it in real-time.
Static. Bayesian optimization is often used to auto-tune parameters for black-box systems [snoek2012practical, shahriari2015taking, frazier2018tutorial]. By iteratively updating its model of the objective function, it efficiently explores the parameter space, tuning planning parameters for better performance.
Dynamic. Dynamic meta-planners adjust parameters on-the-fly for different scenarios [xiao2021learning, Binch2020, xiao2020appld, xu2021applr]. For example, [xiao2020appld] learns a library of parameter sets and switches based on context, while [xu2021applr] uses reinforcement learning to optimize parameter policies over decision sequences.
II-C RL as Meta-Planner
While existing works [gu2022constrainedRL, gehring2022rl4classical] combine RL with rule-based classical planning to enhance global path planning, our paper focuses on integrating RL with a local planner.
We formulate the navigation problem with a meta-planner as a Markov Decision Process (MDP) $\mathcal{M}$, i.e., a tuple $(\mathcal{S},\mathcal{A},\mathcal{T},\gamma,\mathcal{R})$. $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the state transition function, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow R$ is the reward function, and $\gamma$ is the discount factor. At each time step t, the RL agent takes the current state $s_{t}\in\mathcal{S}$ as input and outputs an action $a_{t}\in\mathcal{A}$. Then the environment will transit to the next state $s_{t+1}\sim\mathcal{T}(\cdot|s_{t},a_{t})$ and the agent will receive a reward $r_{t}=\mathcal{R}(s_{t},a_{t})$. The overall objective of RL agent is to learn a policy $\pi:\mathcal{S}\rightarrow A$ that can be used to select actions that maximize the expected cumulative reward over time, i.e. $J=\mathbb{E}_{(s_{t},a_{t})\sim\pi}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}].$
Within the MDP, we optimize a meta-planner policy that interacts with a composite environment consisting of the navigation world $\mathcal{W}$ (filled with obstacles) and a configurable motion planner $f$ parameterized by $\theta$ and receiving sensory inputs $o$. The state space, action space, reward function, and termination conditions are defined as follows:
State space. Following [xu2021applr], we assume a local goal $g$ is always reachable and acts as a waypoint on a global path. The direction to $g$, $\phi\in[-\pi,\pi]$, relative to the robot’s orientation, provides directional information. Thus, the state is $s_{t}=(o_{t},\phi_{t},\theta_{t-1})$, where $o_{t}$ are sensory inputs, $\phi_{t}$ the angle to $g$, and $\theta_{t-1}$ the previous planner configuration.
Action space. The agent’s action $a_{t}=\pi(s_{t})$ updates the planner’s parameters $\theta_{t}$. Executing $a_{t}$ applies $\theta_{t}$ to planner $f$, moving the robot and transitioning to the next state $s_{t+1}$.
Reward. The agent receives an immediate reward $r_{t}$ inversely related to its distance to $g$. An episode ends when a maximum time $T$ is reached or a termination condition is triggered, with the final reward penalizing timeouts and collisions.
The goal of the RL agent within the meta-planner is to learn a policy $\pi$ that can select a planner parameter $\theta_{t}$ for $f$ to achieve optimal navigation performance. The objective can be formulated as:
$\max_{\begin{subarray}{c}\pi\end{subarray}}J^{\pi}=\mathbb{E}_{s_{0},\theta_{t}\sim\pi(s_{t}),s_{t+1}\sim\mathcal{T}(s_{t},\theta_{t})}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}r_{t}\Big{]}.$ (1)
II-D Prioritized Experience Replay
Prioritized Experience Replay [schaul15prioritized] improves RL by assigning higher probabilities to more informative experiences in the replay buffer, enhancing learning efficiency. It selects samples that significantly contribute to the agent’s progress. In contrast, our approach focuses on improving data collection quality by reducing sampling bias and promoting better exploration, resulting in a more informative replay buffer. Together, these methods complement each other to advance RL efficiency.
II-E Roadmap Analysis
Existing roadmap analysis [kavraki1998roadmapanalysis, hsu2006foundationroadmap] often relies on static geometric properties to address challenges like narrow passages, using concepts such as visibility and lookout. While these studies provide valuable insights, they do not fully capture the dynamic execution status resulting from a robot’s interaction with environment. Our approach aims to enhance performance by leveraging robot behavior during dynamic execution. By analyzing the robot’s behaviors as it operates, we identify high-resistance areas that may not be apparent through static analysis alone. This behavior-guided method allows us to adaptively improve planning and navigation, leading to better performance than static geometric analysis.
III Sampling Skew of Existing Methods
APPLR [xu2021applr], the state-of-the-art dynamic meta-planner, exhibits a sampling skew that can lead to inefficiencies and limit performance improvements. In this section, we discuss the sampling skew, key insights, and related technical challenges.
(a) Iteration-1
(b) Iteration-10
(c) Iteration-30
(d) Iteration-90
Figure 3: Sampling skew in existing RL-based Meta-Planner. For convenient visualization, we show 2D projections of the difficult map from Fig. 2 and omit the robot body. The start position is on the left of the map and the goal is on the right. Red cross markers are the robot trajectories at each timestamp. On top and right of the maps are the frequency distribution of robot positions $x$ and $y$.
Meta-Planner Training. Fig. 3 shows 2D projections of the difficult map from Fig. 2, where the robot navigates obstacles from the start (left) to the goal (right) without collisions and within a time limit. The meta-planner adjusts planner’s configurations, i.e. actions $a_{t}=\pi(s_{t})$, at a fixed frequency (e.g., 1 Hz) in response to environmental changes, and we record the robot’s positions for reward calculation and policy training.
Fig. 3(a) shows the early-stage challenges with a randomly initialized policy, where the robot often gets stuck (Area A). Over time (Fig. 3(b)), the policy improves, avoiding early obstacles but still struggling with tight spaces. As training progresses (Fig. 3(c)), the robot reduces getting stuck but takes inefficient detours (Area C), leading to timeout. Further training (Fig. 3(d)) improves navigation, but the robot still faces difficulties passing through narrow gates (Area D).
Sampling Skew Phenomenon. From training iterations in Fig. 3, sampling skew can be summarized by 3 observations:
Observation I: Challenging scenarios require more sampling. In difficult situations like dead corners (Fig. 3(b)), detours (Fig. 3(c)), and narrow gates (Figures 3(a), 3(d)), the classical planner’s limitations and insufficient sampling of the meta-planner often cause delays, timeouts, or collisions. These critical hard-to-pass areas require significantly more sampling to improve navigation efficiency and safety.
Observation II: Challenging scenarios have impact on subsequent scenarios. Hard-to-pass areas make the following scenarios (i.e., hard-to-reach) more difficult to access. This is reflected in the robot’s position distribution, showing higher frequency near hard-to-pass areas (e.g., Area A in Fig. 3(a) and Area B in Fig. 3(b)) and lower frequency in subsequent regions.
Observation III: Challenging scenarios tend to be under-fitted. Certain challenging scenarios remain under-fitted, risking mission success due to two key factors. For instance, Area D initially remains hard-to-reach (Fig. 3(a)-3(c)) and is not effectively trained. As training progresses (Fig. 3(d)), Area D becomes hard-to-pass, as the robot reaches it smoothly but takes around 20 steps (about 20 seconds) per trajectory. This lengthy process limits efficient training within a fixed budget, making it difficult to fully train Area D.
Challenges. A simple solution to mitigate sampling skew is to focus on training the meta-planner more in under-fitted areas, but there is no automatic way to identify these regions. Analyzing static maps alone is insufficient (e.g., detecting narrow parts or sharp turns), as this only considers static information and ignores how the robot dynamically executes and interacts with its environment.
Our key insight is that hard-to-reach areas stem from preceding hard-to-pass areas. If there are no hard-to-pass areas, the robot can easily and safely reach its goal. Therefore, our robot behavior-guided method focuses on: (1) increasing sampling in hard-to-pass areas based on Observations I and II, and (2) reducing the time needed to initiate such sampling, aligning with Observation III. In Section IV, we describe how we identify these areas to address issues in meta-planner.
IV Design of Digimon
We introduce Digimon, consisting of two components: diagnosis (Section IV-B) and mitigation (Section IV-C) of sampling skew phenomenon, as discussed in Section III.
IV-A Overview
Alg. 1 illustrates the integration of Digimon within the meta-planner policy training pipeline. Components enhanced by Digimon are highlighted in blue, indicating the modifications to the standard meta-planner policy. The remaining parts abstract the conventional components of a reinforcement learning training pipeline.
To better quantitatively pinpoint the hard-to-pass areas discussed in Section III, we develop Alg. 2 and define the concept high-resistance area ($\mathcal{H}$), composed of specific poses termed high-resistance points. Intuitively, high-resistance refers to the observation of abrupt behavioral changes by the robot at these points, suggesting that these areas pose significant uncertainty to the robot’s behavior.
Algorithm 1 Digimon for Enhancing RL-based Meta-Planner Training
1:Input: Initial meta-planner policy network $\pi$, high-resistance area $\mathcal{H}$, environment $\mathcal{E}=f\times\mathcal{W}$ for $\pi$ (where $f$ is classical planner and $\mathcal{W}$ is physical map).
2:Constants: Robot’s initial position ${\tt INIT\_POSE}$, threshold $\lambda$ for sampling initial pose from $\mathcal{H}$, policy training iteration $N$, trajectory collection number $K$, policy update number $L$.
3:Output: The converged meta-planner policy network $\pi^{*}$.
4:Set $\mathcal{D}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{P}=\mathcal{H}=\emptyset}$. $\triangleright$ transition buffer $\mathcal{D}$, pose buffer $\mathcal{P}$
5:for iteration $n=0,...,N$ do
6:    for $k=0,...,K$ do
7:        Initialize state $s_{0}$ by resetting $\mathcal{E}$.
8:        Set the robot’s initial pose $p_{0}={\tt INIT\_POSE}$.
9:        Set transition buffer $\mathcal{D}_{k}=\emptyset$, robot pose buffer $\mathcal{P}_{k}=\emptyset$.
10:        if ${\tt getRandom}($0,1)))) < $\lambda$ and $\mathcal{H}\neq\emptyset$ then
11:           $p_{0}\leftarrow{\tt getRandom}(\mathcal{H})$
12:        Set $\tilde{s}=[s_{0}]$ and $\tilde{p}=[p_{0}]$.
13:        for $t=0,...,T$ do
14:           Run $\pi$ in $\mathcal{E}$, take action $a_{t}=\pi(s_{t})$.
15:           Obtain reward $r_{t}$, state $s_{t+1}$ and pose $p_{t+1}$.
16:           $\tilde{s}\leftarrow\tilde{s}.{\tt append}(s_{t+1})$,   $\tilde{p}\leftarrow\tilde{p}.{\tt append}(p_{t+1})$.
17:           if $\pi$ gets ${\tt done}$ signal from $\mathcal{E}$ then,
18:               $\mathcal{D}_{k}\leftarrow\mathcal{D}_{k}\cup\{\tilde{s}\}$, $\mathcal{P}_{k}\leftarrow\mathcal{P}_{k}\cup\{\tilde{p}\}$.
19:        Update $\mathcal{D}$ using $\mathcal{D}_{k}$. $\triangleright$ This step is algorithm-specific
20:        Update $\mathcal{H}$: $\mathcal{H}\leftarrow{\tt getHRArea}(\mathcal{P}_{k})$. $\triangleright$ Refer to Alg. 2.
21:    for $\ell=0,...,L$ do
22:        $\mathcal{B}\leftarrow{\tt getMiniBatch}(\mathcal{D})$triangleright Random mini-batch
23:        Update policy $\pi$ using $\mathcal{B}$. $\triangleright$ Algorithm-specific
24:Return the final meta-planner policy $\pi^{*}$.
IV-B Sampling Skew Diagnosis
High-Resistance Point Identification. Alg. 2 identifies high-resistance areas $\mathcal{H}$ within robot’s navigation environment. By processing pose trajectories $\mathcal{P}$ and a threshold $\eta$, it outputs $\mathcal{H}$, i.e. poses where the robot exhibits significant behavioral changes, indicating challenging regions for the planner.
The algorithm begins by accepting a set of pose trajectories $\mathcal{P}$. For each trajectory in $\mathcal{P}$, the algorithm checks if the trajectory successfully reaches the mission goal, filtering out any mission-failure trajectories (in Line 5 of Alg. 2). For trajectories that reach the mission goal, the algorithm examines each triplet of consecutive poses $(p_{i},p_{i+1},p_{i+2})$ within the trajectory. The poses are characterized by their position $x,y$ and orientation $w$.
In Line 11 of Alg. 2, the differences in $x$ and $y$ coordinates between consecutive robot poses are used to calculate trajectory orientations, notated as $\rho_{i}=\arctan(\Delta y_{i}/\Delta x_{i})$ and $\rho_{i+1}=\arctan(\Delta y_{i+1}/\Delta x_{i+1})$. Then variation of trajectory orientations $\Delta\rho_{i}=\rho_{i+1}-\rho_{i}$ is then normalized to the range $[-\pi,\pi]$ (in Line 12 of Alg. 2). If the absolute value $|\Delta\rho_{i}|$ exceeds the threshold $\eta$, the corresponding pose $p_{i}$ is identified as a high-resistance point and added to the set $\mathcal{H}$. Finally, the algorithm returns the set $\mathcal{H}$, representing the high-resistance area. The key intuition behind Alg. 2 is to capture regions where the meta planner shows high uncertainty in challenging scenarios, indicated by a high variance in the trajectory.
Note that there are two technical aspects that merit particular attention: ❶ Alg. 2 employs variations in trajectory orientations $\Delta\rho$ rather than robot orientations $\Delta w$ as the metric for comparison against the threshold $\eta$. This design effectively captures behaviors such as back-and-forth movements, where the robot’s orientation may remain relatively unchanged and thus cannot be detected by $\Delta w$. However, these movements are also clear indicators of challenges faced by the planner. ❷ Alg. 2 strategically excludes failure trajectories. This exclusion ensures that any points of high resistance can ultimately reach the goal. Allowing the robot to persist in areas that are excessively challenging would cause the planner to become indefinitely stuck, thereby introducing noise into the training of the meta-planner. This process allows for the identification of challenging areas where the robot experiences significant difficulties, enabling targeted interventions to improve the planner’s performance in these regions.
Algorithm 2 High-Resistance Area Identification ${\tt getHRArea}$
1:Input: Pose trajectory set $\mathcal{P}$, threshold $\eta$ for high-resistance identification.
2:Output: High-resistance area $\mathcal{H}$.
3:Initialize $\mathcal{H}=\emptyset$.
4:for each $\tilde{p}$ in $\mathcal{P}$ do
5:    if $\tilde{p}$ reaches the mission goal then
6:        $\triangleright$ $p$ is a 3-dim vector: position $x,y$ and orientation $w$
7:        for each $(p_{i},p_{i+1},p_{i+2})$ in $\tilde{p}$ do
8:           $(x_{k},y_{k},w_{k})\leftarrow p_{k},k=\{i,i+1,i+2\}$, .
9:           $\Delta x_{i},\Delta x_{i+1}\leftarrow x_{i+1}-x_{i},x_{i+2}-x_{i+1}$.
10:           $\Delta y_{i},\Delta y_{i+1}\ \leftarrow y_{i+1}-y_{i},y_{i+2}-y_{i+1}$.
11:           $\rho_{i},\rho_{i+1}\leftarrow\arctan\Big{(}\dfrac{\Delta y_{i}}{\Delta x_{i}}\Big{)},\arctan\Big{(}\dfrac{\Delta y_{i+1}}{\Delta x_{i+1}}\Big{)}$
12:           $\Delta\rho_{i}\leftarrow\rho_{i+1}-\rho_{i}$.
13:           $\Delta\rho_{i}\leftarrow(\Delta\rho_{i}\mod 2\pi)-\pi$. $\triangleright$ Normalization
14:           if $|\Delta\rho_{i}|>\eta$ then $\mathcal{H}\leftarrow\mathcal{H}\cup\{p_{i}\}$
15:Return the high-resistance area $\mathcal{H}$.
IV-C Sampling Skew Mitigation
High-Resistance Area Up-sampling. In Lines 10-11 of Alg. 1, the process of high-resistance area up-sampling is implemented to enhance the robustness and adaptability of the meta-planner during training. This algorithm leverages threshold $\lambda\in[0,1]$ to determine whether to sample the robot’s initial position from the set of high-resistance areas $\mathcal{H}$. It ensures that the meta-planner frequently encounters and learns to navigate through high-resistance regions, thereby improving its ability to handle challenging scenarios. By balancing this targeted exploration with a general exploration of the environment, the meta-planner will train more efficiently and effectively in high-resistance areas. Additionally, $\mathcal{H}$ is updated iteratively in Line 20 of Alg. 1, ensuring that the set of high-resistance areas remains relevant and accurately reflects the current sampling challenges. This iterative update is crucial for maintaining the effectiveness of the up-sampling strategy. Overall, the up-sampling of high-resistance areas is critical for the meta-planner’s training as it enhances the planner’s robustness, promotes better generalization to new environments, and addresses sampling skew.
V Evaluation
Table I: Navigation performance of Digimon and baseline methods under Same-Env setup.
Metric $\Rightarrow$ Method $\Downarrow$ Easy Medium Difficult
NS $\uparrow$ ATT $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$ NS $\uparrow$ ATT $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$ NS $\uparrow$ ATT $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$
DWA 20.72 30.01 97.65 1.18 1.18 16.91 33.35 90.59 2.35 7.06 10.61 30.99 53.75 16.25 30.00
DWA-Fast 25.44 23.60 76.47 23.53 0.00 16.30 30.68 64.71 23.53 11.76 12.43 27.40 50.00 18.75 31.25
BayesOpt 17.88 48.09 74.00 10.00 16.00 15.61 21.84 49.02 21.57 29.41 5.94 58.37 43.75 6.25 50.00
APPLR 20.19 28.48 70.59 13.24 16.18 21.36 26.36 85.29 7.35 7.35 12.83 33.59 57.81 4.69 37.50
Ours 35.03 13.45 80.88 11.76 7.35 23.23 21.62 76.74 17.65 5.88 14.56 27.81 60.94 18.75 20.31
  • NS: Navigation Score. ATT: Actual Traveling Time. SR: Success Rate. CR: Collision Rate. TR: Timeout Rate.
Table II: Navigation performance of Digimon and baselines under Cross-Env and Cross-Level setups. The column names represent the test environment levels: All (across all three levels), Easy, Medium and Difficult.
Training Env. Metric $\Rightarrow$ Method $\Downarrow$ All Easy Medium Difficult
NS $\uparrow$ ATT $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$ NS $\uparrow$ ATT $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$ NS $\uparrow$ ATT $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$ NS $\uparrow$ ATT $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$
Easy BayesOpt 12.67 44.58 72.00 10.00 18.00 18.33 37.00 88.24 0.00 11.76 12.68 45.81 76.47 17.65 5.88 6.65 56.79 50.00 12.50 37.50
APPLR 10.46 37.91 47.60 19.60 32.80 20.04 29.82 71.76 12.94 15.29 7.06 44.63 44.71 17.65 37.65 3.90 49.80 25.00 28.75 46.25
Ours 20.43 19.04 58.40 15.20 26.40 35.32 14.08 82.35 11.76 5.88 15.26 22.16 50.59 22.35 27.06 10.10 25.48 41.25 11.25 47.50
Medium BayesOpt 19.74 17.68 52.00 16.00 32.00 35.18 10.78 76.47 5.88 17.65 13.38 25.33 41.18 29.41 29.41 10.09 23.72 37.50 12.50 50.00
APPLR 21.85 25.39 77.60 11.60 11.80 31.43 20.71 90.59 9.41 0.00 21.49 26.89 84.71 7.06 8.24 12.06 31.02 56.25 18.75 25.00
Ours 24.39 22.22 75.20 13.60 11.20 37.14 15.66 92.94 7.06 0.00 22.61 22.40 75.29 18.82 5.88 12.73 33.47 56.25 15.00 28.75
Difficult BayesOpt 8.72 52.23 52.00 10.00 38.00 11.32 43.11 58.82 5.88 35.29 8.87 54.66 52.94 5.88 41.18 5.81 62.13 43.75 18.75 37.50
APPLR 22.56 25.21 76.00 8.40 15.60 34.51 18.93 92.94 5.88 1.18 19.65 26.69 75.29 12.94 11.76 12.97 33.74 58.75 6.25 35.00
Ours 24.61 22.54 79.60 12.00 8.40 36.30 17.90 96.47 3.53 0.00 22.16 24.53 80.00 12.94 7.06 14.78 27.56 61.25 20.00 18.75
  • NS: Navigation Score. ATT: Actual Traveling Time. SR: Success Rate. CR: Collision Rate. TR: Timeout Rate.
V-A Experiment Setup
Simulation & Planners. Following existing works [xiao2020appld, xu2021applr], we implement Digimon using the widely-used Gazebo [Gazebo] simulator with a ClearPath Jackal differential drive ground robot as a robot model. Due to space limit, we leave the details of the classical planner and the meta-planner that Digimon builds upon for evaluation to Appendix VII.
Baselines & Dataset. We compare Digimon with two widely used classical planners (DWA [fox1997dwa] and DWA-Fast [dwafast]), one representative static meta-planner, BayesOpt [snoek2012practical] and one SOTA dynamic RL-based meta-planner APPLR [xu2021applr]. We choose the same dataset as APPLR, namely BARN [perille2020benchmarking] dataset, as it includes various maps for robot navigation inside the Gazebo simulator and it has different difficulty levels’ map (i.e., Easy, Medium, and Difficult) to evaluate. This allows for a convenient evaluation of various planners’ performance across these varied settings.
Metrics. We evaluate navigation systems using five metrics: ❶ Actual Traveling Time (ATT): Time (in seconds) to complete successful missions. ❷ Mission Success Rate (SR): Percentage of missions completed (i.e., reaching the goal without collisions and within the time budget.). ❸ Collision Rate (CR): Percentage of missions failed due to collisions with obstacles. ❹ Timeout Rate (TR): Percentage of missions where the robot failed to reach the goal within the time limit. ❺ Navigation Score (NS): A comprehensive performance metric from BARN Challenge [barn_challenge], calculated as follows:
$\small{\rm{\bf NS}_{i}}=\mathbb{I}({\text{Success}})\times\dfrac{OT_{i}}{{\rm clip}({\bf ATT}_{i},2OT_{i},8OT_{i})}$ (2)
Here, $OT_{i}$ represents the optimal time (computed using Dijkstra algorithm [dijkstra2022note] in advance), and ${\rm clip}({\bf ATT},2OT_{i},8OT_{i})$ ensures that the actual traveling time is constrained within twice and eight times the optimal time. $\mathbb{I}(\textrm{Success})$ is an indicator function that denotes whether the current navigation mission is successful, consistent with the definition of SR. Higher values of NS and SR indicate better performance, while lower values of ATT, CR, and TR indicate better performance.
Comparison Setup. We introduce 3 evaluation setups: (1) Same-Env: The meta-planner is trained and tested in the same environment set with predefined but different start positions to avoid overfitting. (2) Cross-Env: The meta-planner is trained and tested at the same difficulty level but in different environments. (3) Cross-Level: The meta-planner is trained on one difficulty level but tested on a different difficulty level.
V-B Same-environment Performance
We evaluate Digimon and four baselines under the Same-Env setup. DWA, DWA-Fast, and BayesOpt are static planners, so we directly implement their fixed configurations from the training environment and test them on the same environments with random starting positions. For APPLR and Digimon, we train the meta-planner policy on the training set and deploy it on the test environments.
Figure 4: Navigation score (NS) of different checkpoints.
Table III: Performance of Digimon on PPO-based meta-planner.
PPO + APPLR PPO + Ours
NS $\uparrow$ 17.58 20.61
ATT $\downarrow$ 24.69 22.07
SR $\uparrow$ 55.56 65.28
CR $\downarrow$ 26.39 15.28
TR $\downarrow$ 18.06 19.44
Table IV: Performance of Digimon w/ and wo/ failure filtering.
wo/ Filtering w/ Filtering
NS $\uparrow$ 22.22 23.23
ATT $\downarrow$ 19.60 21.62
SR $\uparrow$ 70.59 76.47
CR $\downarrow$ 22.06 17.65
TR $\downarrow$ 7.35 5.88
Table I shows that Digimon consistently achieves the highest NS score across all difficulty levels, demonstrating its ability to identify and improve performance in underrepresented scenarios. While DWA performs better on SR and CR in easy and medium environments, Digimon excels in the comprehensive NS metric. DWA’s conservative design prioritizes safety over time efficiency, leading to longer navigation times, as seen in reduced ATT and lower NS scores, making it less practical for navigation tasks. Since NS is a widely-used metric [xu2021applr, xiao2020appld], Digimon’s higher NS score effectively demonstrates its superiority over both static planners and the SOTA dynamic meta-planner.
V-C Cross-environment & Cross-level Performance
In this section, we compare Digimon and selected baselines under the Cross-Env and Cross-Level setup, as shown in the diagonal sub-tables in Table II. Under the Cross-Env setup, Digimon consistently achieves the highest NS score across all three environments, indicating its effectiveness in identifying and addressing underrepresented scenarios to enhance planner performance. For the Cross-Level experiments, Digimon excels in generalization by achieving the highest NS scores when the trained meta-planner is tested across different environmental difficulty levels. This demonstrates Digimon’s robustness and capability to maintain navigation efficacy in new and varying conditions. Additionally, when meta-planners trained in difficult environments are applied to easier ones, Digimon outperforms even those baselines specifically trained for simpler environments, while this trend is not observed in the baseline methods. This result suggests that the meta-planners trained by our method can learn more sophisticated and adaptable configurations, enabling Digimon to excel in environments that are less challenging.
V-D Diagnosis Effectiveness
To better understand why our diagnosis and mitigation strategy is effective, we introduce a variant called RandomSampling (RS). Unlike Digimon, which identifies high-resistance areas with Alg. 2, RS indiscriminately samples from all locations reached by the robot during training without strategic guidance, while all other designs remain consistent with Digimon. Results in Fig. 5 show that Digimon significantly outperforms RS, with higher NS scores in all environments. The reasons are two-fold. First, RS fails to expose the robot to challenging, hard-to-reach scenarios, limiting effective training in these high-resistance areas, which is a crucial aspect for satisfactory navigation performance. Second, as shown in Fig. 2, medium and hard environments have more complicated maps with many hard-to-reach scenarios and obstacles than the easy level. Random sampling in such environments often leads to the robot encountering “dead-corner” scenarios – areas that are hard to exit, leading to extended, unrewarding training trajectories. Such trajectories will destabilize the training of the RL policy in the meta-planner and degrade navigation performance. This analysis verifies the necessity of our high-resistance area up-sampling step in Digimon. By strategically selecting those hard-to-reach areas, we enhance the robot’s ability to perform well across a variety of navigational scenarios. Appendix VIII shows more results in the comparison between RS and Digimon.
(a) Easy
(b) Medium
(c) Difficult
Figure 5: Comparison of RandomSampling and Digimon. The meta-planners are trained on three different difficulty levels and tested across all levels.
V-E Efficiency
In this section, we evaluate the performance of different meta-planner policy checkpoints during training under the Cross-Env setup when Digimon is applied in a medium-level environment. Fig. 4 shows the average NS metric when those checkpoints are evaluated in testing environments. Digimon consistently outperforms APPLR, achieving higher navigation scores at all training stages with the same training steps. Furthermore, APPLR requires over four times training steps to achieve comparable performance of Digimon, and our method ultimately outperforms APPLR by 13.5%. These results demonstrate that our method significantly enhances the training efficiency of the RL policy within the meta-planner and improves the overall navigation performance of the robot. Besides NS, we also show more metrics of different meta-planner checkpoints during training (in Fig. 9 in Appendix).
V-F Plug-and-Play Applicability on Other RL Algorithm
In Sections V-B and V-C, we have demonstrated that Digimon enhances navigation performance when paired with the off-policy algorithm TD3 [fujimoto2018addressing] based meta-planner in APPLR. To further validate the versatility of Digimon, we extended it to another SOTA on-policy RL algorithm, Proximal Policy Optimization (PPO) [ppo]. For a fair comparison, we replace the TD3 training algorithm in APPLR with PPO. Results are shown in Table V-B. We can observe that Digimon outperforms baseline methods when applied to PPO, where four out of five metrics get significantly improved, and a negligible drop in timeout rate. This indicates that compared to APPLR, our diagnostic method is robust across various training algorithms and not limited to a specific algorithm.
V-G Other Experiments
Ablation study. We conduct the ablation study to justify the importance of key components in Digimon and the sensitivity test against the key hyper-parameters. Details can be found in Appendix IX-B.
VI Conclusion
In this paper, we identify and analyze the challenges of integrating RL with classical planning in meta-planners. While this integration enhances adaptability and explainability, it introduces sampling skew due to the conservative nature of classical planners, limiting the RL meta-planners’ exploration and performance in complex environments.
To address these issues, we propose Digimon, a diagnostic and mitigation framework for RL-based meta-planners. Digimon analyzes robot behaviors, identifies exploration bottlenecks and applies a targeted up-sampling strategy to balance the sampling data, enabling more effective learning and exploration. Our experiments demonstrate that Digimon improves navigation performance by 13.5%+, enhances robustness in unseen environments, and quadruples training efficiency.
Limitations. Although Digimon greatly boosts efficiency, further optimization of training duration is possible. The current pipeline depends on executing the classical planner to collect data, meaning increased complexity of the classical planner can raise training costs. Additionally, Digimon is specifically designed for navigation tasks and may not apply to manipulation tasks, which we leave for future work.
Appendix
VII Experiment Setup
We run experiments on Ubuntu 20.04, with Intel i9-13900K, 64 GB RAM and Nvidia GPU RTX 2080. The robot is equipped with a 720-dimensional planar laser scanner with a 270 field of view, which provides our sensory input $o_{t}$. We pre-process the LiDAR scans by capping the maximum range to 2m. The onboard Robot Operating System (ROS) ${\tt move\_base}$ navigation stack (our underlying classical motion planner $f$) uses Dijkstra’s algorithm [dijkstra2022note] to plan a global path and runs DWA [fox1997dwa] as the local classical planner to follow the global path.
The meta-planner policy $\pi$ is trained to update the planning configurations $\theta$ of the classical planner. We maintain the same settings as APPLR, utilizing the DWA classical planner and the same action space $\theta$, including ${\tt max\_vel\_x}$, ${\tt max\_vel\_theta}$, ${\tt vx\_samples}$, ${\tt vtheta\_samples}$, ${\tt occdist\_scale}$, ${\tt pdist\_scale}$, ${\tt gdist\_scale}$, and ${\tt inflation\_radius}$ [^1] https://wiki.ros.org/dwa_local_planner We use the ROS dynamic reconfigure client to dynamically change planner parameters. The global goal $\alpha$ is assigned manually, while $\theta_{0}$ is the default set of parameters provided by the robot manufacturer. [^2] http://wiki.ros.org/costmap_2d Table V shows the details of each configuration of the meta-planner.
Configuration Min Max Meaning
${\tt max\_vel\_x}$ 0.1 2.0 m/s. Maximum threshold for linear velocity.
${\tt max\_vel\_theta}$ 0.314 3.14 rad/s. The absolute value of the maximum rotational velocity for the robot.
${\tt vx\_samples}$ 4 12 The number of samples to use when exploring the x velocity space.
${\tt vtheta\_samples}$ 8 40 The number of samples to use when exploring the theta velocity space
${\tt path\_distance\_bias}$ 0.1 0.5 The weight for how much the robot should stay close to the path it was given.
${\tt goal\_distance\_bias}$ 0.1 2 The weight for how much the robot should attempt to reach its local goal.
${\tt inflation\_radius}$ 0.1 0.6 Controls how far away the zero cost point is from obstacle.
Table V: Action Space of Meta-Planner Policy $\pi$.
VIII Diagnosis Effectiveness
In this section, we delve deeper into the comparative analysis between RandomSampling and Digimon by applying both methods across environments of varying complexity: Easy, Medium, and Difficult. We conducted separate experiments for each difficulty level, which allowed for detailed and direct comparisons. The results of these experiments are represented in Figures 67, and 8.
Consistent with the observations discussed in Section V-D, our Digimon continues to outperform RandomSampling in overall navigation performance. This superiority is quantified by a higher Navigation Score (NS), indicating Digimon’s effectiveness in handling complex navigation tasks. This comparison not only highlights the strengths of Digimon but also contributes to a better understanding of its operational dynamics across different environmental settings.
(a) Train on Easy
(b) Train on Medium
(c) Train on Difficult
Figure 6: Test on Easy Level
(a) Train on Easy
(b) Train on Medium
(c) Train on Difficult
Figure 7: Test on Medium Level
(a) Train on Easy
(b) Train on Medium
(c) Train on Difficult
Figure 8: Test on Difficult Level
Table VI: Navigation performance under different high-resistance area up-sampling ratios $\lambda$.
Ratio $\lambda$ NS $\uparrow$ ATT $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$
0.2 9.63 27.01 38.75 25.00 36.25
0.4 13.33 31.63 57.50 12.50 30.00
0.6 11.18 38.38 55.00 13.75 31.25
0.8 11.59 27.04 42.50 17.50 40.00
Table VII: Navigation performance under different trajectory vector change thresholds $\eta$.
Threshold $\eta$ () NS $\uparrow$ ATT $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$
50 15.19 26.89 50.59 18.82 30.59
70 18.92 26.78 65.20 14.80 20.00
90 22.61 22.40 75.29 18.82 5.88
110 18.22 28.93 74.12 15.29 10.59
130 17.45 23.70 57.65 14.12 28.24
(a) Average Time
[…]
(b) Average Reward
[…]
(c) Average Success Rate
Figure 9: Average time, reward, and success rate of Digimon and APPLR for different checkpoints during training.
IX Additional Experiments
IX-A More Results of Efficiency Comparison
In Figure 9, we present additional metrics for various meta-planner checkpoints during training, including average traveling time, reward, and success rate. It clearly shows that Digimon outperforms APPLR across all three metrics. Notably, Digimon requires fewer training steps to achieve shorter traveling times, highlighting the method’s efficiency. Similarly, within the same amount of training steps, Digimon can also achieve a higher average reward and success rate. All of these three metrics robustly demonstrate that Digimon can achieve superior efficacy compared to the SOTA baseline APPLR.
IX-B Ablation Study
Sensitivity. We conduct ablation studies to evaluate the sensitivity of Digimon against various crucial hyper-parameters: (1) high-resistance area sampling threshold $\lambda$ in Algorithm 1 and (2) high-resistance point identification threshold $\eta$ in Algorithm 2.
Failure Trajectory Filtering. We evaluate the necessity of our failure trajectory filtering in Digimon and its advantages for improving navigation performance. The results are shown in Table V-B. We observe that without the failure filtering trajectory mechanism, three out of five metrics deteriorate significantly, particularly the collision rate and success rate, which are critical to the robot’s safety. In contrast, without failure filtering, the NS and ATT metrics show only marginal improvements. This indicates that filtering out failure trajectories can substantially enhance overall navigation performance in terms of safety and task completion rate while not sacrificing total travel time.
High-resistance Area Up-sampling Ratio $\lambda$. We then study the influence of $\lambda$ in Alg. 1 on the navigation performance. Table VIII shows that $\lambda=0.4$ yields the best performance for the navigation system. Varying $\lambda$ introduces fluctuations across various metrics, while all within a 10% range. This demonstrates the necessity of fine-tuning the high-resistance area up-sampling ratio $\lambda$ to balance the proportion of difficult trajectories within the total training set. If $\lambda$ is set too high, the training set will contain too many challenging trajectories, complicating the meta-planner’s ability to understand the complete map and complete the task. If $\lambda$ is set too low, it will revert to the original training method, hindering the meta-planner’s ability to learn from those under-fitted trajectories.
In Table VIII, we show the performance metrics of Digimon over different threshold angles. While there are some improvements in certain metrics (e.g., NS and CR) at higher threshold angles, the performance data clearly converges around 90° as the optimal threshold for overall system performance. This peak suggests at 90°, the system achieves a balance between speed, efficiency, and safety, highlighting a degree of robustness and adaptability. Similar to the high-resistance area up-sampling ratio $\lambda$, the threshold angle should also be carefully tuned to enable the best performance. Thus, we set $\lambda=0.4$ and $\eta=90°$ for all other experiments.
X Potential Social Impacts
The development and deployment of Digimon, aimed at enhancing the performance of meta-planner policy in robotic navigation, promises significant positive social benefits. By enabling robots to navigate more effectively and reliably, our tool can greatly facilitate the integration of robotics into our daily life. For example, in healthcare settings, more efficient navigation can enable robots to deliver medication, assist in surgeries, or provide companionship with greater precision and safety. In homes, robots equipped with advanced navigation capabilities can assist individuals with disabilities, offering them greater independence and quality of life. Furthermore, in disaster response situations, robots that can navigate challenging terrains could save lives by reaching areas that are inaccessible to humans. Overall, by improving the reliability and functionality of meta-planners in robotic navigation systems, Digimon has the potential to make significant contributions to society, improving safety, accessibility, and efficiency in numerous critical areas.

### v3

@@
 Abstract
-Robot navigation is increasingly crucial across applications like delivery services and warehouse management.
-The integration of Reinforcement Learning (RL) with classical planning has given rise to meta-planners that combine the adaptability of RL with the explainable decision-making of classical planners.
-However, the exploration capabilities of RL based meta-planners during training are often constrained by the capabilities of the underlying classical planners.
-This constraint can result in limited exploration, thereby leading to sampling skew issues.
-To address these issues, our paper introduces a novel framework, Digimon, which begins with behavior-guided diagnosis for exploration bottlenecks within the meta-planner and follows up with a mitigation strategy that conducts up-sampling from diagnosed bottleneck data.
-Our evaluation shows 13.5%+ improvement in navigation performance, greater robustness in out-of-distribution environments, and a 4$\times$ boost in training efficiency.
-Digimon is designed as a versatile, plug-and-play solution, allowing seamless integration into various RL-based meta-planners.
+Robot navigation systems are critical for various real-world applications such as delivery services, hospital logistics, and warehouse management.
+Although classical navigation methods provide interpretability, they rely heavily on expert manual tuning, limiting their adaptability.
+Conversely, purely learning-based methods offer adaptability but often lead to instability and erratic robot behaviors.
+Recently introduced parameter tuners aim to balance these approaches by integrating data-driven adaptability into classical navigation frameworks.
+However, the parameter tuning process currently suffers from training inefficiencies and redundant sampling, with critical regions in environment often underrepresented in training data.
+In this paper, we propose EffiTune, a novel framework designed to diagnose and mitigate training inefficiency for parameter tuners in robot navigation systems.
+EffiTune first performs robot-behavior-guided diagnostics to pinpoint critical bottlenecks and underrepresented regions.
+It then employs a targeted up-sampling strategy to enrich the training dataset with critical samples, significantly reducing redundancy and enhancing training efficiency.
+Our comprehensive evaluation demonstrates that EffiTune achieves more than a 13.5% improvement in navigation performance, enhanced robustness in out-of-distribution scenarios, and a 4× improvement in training efficiency within the same computational budget.
 I Introduction
@@
 Navigating cluttered environments remains a key challenge for achieving agile, efficient, and safe movement.
-Classical planners [abdulsaheb2022robot, asano1985visibility, Canny1985AVM, Sampling2014] address this by using specific cost functions and rules, which are difficult to generalize and often require expert manual tuning [Campbell2020, raheem2022robot, zheng2021rostuning, feng2024rocas].
-In contrast, learning-based approaches [pfeiffer2017perception, gao2017intention, chiang2019learning, xiao2021toward, liu2021lifelong, xiao2021agile, xiao2020appld] offer adaptability but can suffer from instability and lack interpretability, leading to erratic behaviors that compromise safety and efficiency.
-To overcome these limitations, recent studies introduced meta-planners [xu2021applr, xiao2022motion] that integrate RL with classical planning frameworks (see Fig.
+Classical navigation systems [abdulsaheb2022robot, asano1985visibility, Canny1985AVM, Sampling2014] address this by using specific cost functions and rules, which are difficult to generalize and often require expert manual tuning [Campbell2020, raheem2022robot, zheng2021rostuning, feng2024rocas, kate2025roscallbax].
+In contrast, learning-based approaches [pfeiffer2017perception, gao2017intention, chiang2019learning, xiao2021toward, liu2021lifelong, xiao2021agile, xiao2020appld] offer adaptability but can suffer from instability and lack interpretability [feng2023decree, chen2024temporalrl, cheng2024fusion, cheng2024badpart], leading to erratic behaviors that compromise safety and efficiency.
+To overcome these limitations, recent studies introduced parameter tuner [xu2021applr, xiao2022motion, jiang2022efficient, jiang2023learning] that integrate data-driven tuner with classical planning frameworks (see Fig.
 1).
-RL-based meta-planners dynamically adjust the parameters of classical planners, combining RL’s adaptability with the explainable decision-making of classical methods for balanced robot navigation.
-However, existing meta-planners suffer from sampling skew due to the integration of RL and classical planners.
-The exploration capabilities of RL are often limited by the classical planner’s inherent strategies.
-For instance, if the classical planner prefers conservative paths, the RL meta-planner may be unable to explore more aggressive or unconventional routes, resulting in sub-optimal performance.
-Previous approaches [swazinna2021overcoming, shen2023loose, mark2024offline, Strehl2005ATA, nikishin2022primacy], such as adding noise to RL actions, do not address these specific integration challenges.
-Therefore, there is a crucial need for specialized methods to mitigate the biases introduced by combining RL meta-planners with classical planning.
+These data-driven parameter tuners adjust the parameters (across perception, planning and control components) of classical navigation systems in different scenarios, combining adaptability of data-driven methods with the explainable decision-making of classical methods for balanced robot navigation.
+However, existing parameter tuners encounter significant challenges.
+The interaction between classical navigation systems and the environment is computationally expensive, often leading to sampling inefficiencies.
+Furthermore, the training process of these tuners tends to include redundant sampling.
+In addition, our analysis reveals that only a few key regions in the environment are critical for achieving optimal performance.
+Unfortunately, these pivotal areas are underrepresented in the training data due to both the high cost of environmental interaction and the prevalence of redundant sampling.
+To address these challenges, our approach emphasizes the most informative data during training, significantly enhancing the effectiveness and efficiency of the parameter tuner.
+Unlike static roadmap analysis, which offers a fixed representation, our method captures the robot’s dynamic behavior, providing a more accurate and informative analysis.
+In this paper, we propose EffiTune, a diagnosis and mitigation framework that addresses inefficiency in parameter tuner training.
+EffiTune begins with a robot-behavior-guided diagnostic phase that identifies critical bottlenecks.
+It then employs a novel up-sampling strategy to selectively enrich the training data with samples from these key regions, providing the parameter tuner with a more balanced and informative dataset while reducing redundancy.
+This targeted approach accelerates the training process of tuners, enhances adaptability, and ultimately improves overall navigation performance.
 (a) Classical
-(b) RL End-to-End
-(c) Hybrid
-Figure 1: Different Types of Planners. Note that the meta-planner is applied onto classical local planner $f$, not global trajectory planner.
-In this paper, we propose Digimon, a diagnosis and mitigation framework addressing the limitations of RL meta-planners.
-Digimon starts with a behavior-guided diagnostic phase to identify exploration bottlenecks caused by classical planners.
-It then employs a novel up-sampling strategy to sample data from these bottleneck areas, providing RL meta-planner with a balanced replay buffer.
-This enhances RL exploration ability and mitigates sampling skew caused by classical planner’s preference.
+(b) Hybrid
+Figure 1: Different Types of Robot Navigation Systems. Parameter tuners are typically data-driven and requires training.
 In summary, our contributions are:
-  • We identify and analyze the exploration bottlenecks and sampling skews arising from integrating RL with classical planning in meta-planners.
-  • We present Digimon, featuring a diagnostic phase to detect bottlenecks and a mitigation phase that uses data up-sampling to improve RL meta-planner training.
-  • Digimon achieves 13.5%+ improvement in navigation performance, greater robustness in out-of-distribution environments, and a 4$\times$ increase in training efficiency.
-Additionally, Digimon is a versatile, plug-and-play solution, compatible with various RL-based meta-planners.
+  • We identify training inefficiency issue for parameter tuner training in robot navigation.
+  • We introduce EffiTune, a framework that features a diagnostic phase to detect these critical regions and a mitigation phase that uses targeted data up-sampling to accelerate parameter tuner training.
+  • EffiTune achieves a 13.5%+ improvement in navigation performance, greater robustness in out-of-distribution environments, and a 4× increase in training efficiency within the same timing budget.
 II Background, Notations and Related Work
@@
 2).
-It is worth noting that there are 2 major categories of planning algorithms: local trajectory planners (e.g.
-DWA [fox1997dwa], TEB [teb]) and global path planners (e.g., Dijkstras [dijkstra], A* [astar], RRT [lavalle1998rrt, lavalle2001rrt]).
+There are 2 major categories of planning algorithms: local trajectory planners[fox1997dwa, teb] and global path planners [dijkstra, astar, lavalle1998rrt, lavalle2001rrt].
 In this paper, we consider motion planning in navigation tasks and assume that the robot employs a classical local motion planner $f$.
@@
 The local planner $f$ is responsible for computing the motion commands $u_{t}=f(o_{t},g_{t}|\theta)$ (e.g., angular or linear velocity) at each time step to reach $g_{t}$.
-II-B Meta-Planner
+(a) 0 $\sim$ 2.5 s/m
+(b) 2.5 $\sim$ 4.0 s/m
+(c) 4.0 $\sim$ 8.0 s/m
 Figure 2: Maps of various difficulty levels.
-Meta-Planner operates on top of the original planner $f$, adjusting its configuration $\theta$.
+We use the difficulty metric of “normalized traversal time” (traversal time averaged over the ten trials and normalized by path length) from [perille2020benchmarking] to partition map datasets to 3 levels.
+II-B Parameter Tuner
+Parameter Tuner $\tau$ operates on top of the original planner $f$, taking environmental states $s$ as input and adjusting navigation configuration $\theta$.
 Fig.
-1c shows that a hybrid planner with a meta-planner combines the explainability of classical planners with the adaptability of RL-based planners, enhancing performance.
-It has been applied across domains such as robot navigation [fridovich2018planning, xiao2022motion, siva2019robot, pokle2019deep], autonomous driving [jiang2022efficient], drone control [loquercio2022autotune], and robotic manipulation [Moll2021HyperPlanAF, jia2023dynamic].
-There are two types of meta-planners: static and dynamic. Static meta-planners tune the configuration $\theta$ before deployment, while dynamic meta-planners adjust it in real-time.
+1c shows that a hybrid planner with a parameter tuner combines the explainability of classical planners with the adaptability of data-driven (e.g., end-to-end) planners, enhancing performance.
+It has been applied across domains such as robot navigation [fridovich2018planning, xiao2022motion, siva2019robot, pokle2019deep], self-driving [jiang2022efficient], autonomous system [feng2025intentest, zheng2025llm], drone control [loquercio2022autotune], and robotic manipulation [Moll2021HyperPlanAF, jia2023dynamic].
+There are two types of parameter tuners: static and dynamic. Static parameter tuners tune the configuration $\theta$ before deployment, while dynamic parameter tuners adjust it in real-time.
 Static.
@@
 Dynamic.
-Dynamic meta-planners adjust parameters on-the-fly for different scenarios [xiao2021learning, Binch2020, xiao2020appld, xu2021applr].
-For example, [xiao2020appld] learns a library of parameter sets and switches based on context, while [xu2021applr] uses reinforcement learning to optimize parameter policies over decision sequences.
-II-C RL as Meta-Planner
-While existing works [gu2022constrainedRL, gehring2022rl4classical] combine RL with rule-based classical planning to enhance global path planning, our paper focuses on integrating RL with a local planner.
-We formulate the navigation problem with a meta-planner as a Markov Decision Process (MDP) $\mathcal{M}$, i.e., a tuple $(\mathcal{S},\mathcal{A},\mathcal{T},\gamma,\mathcal{R})$.
-$\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the state transition function, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow R$ is the reward function, and $\gamma$ is the discount factor.
-At each time step t, the RL agent takes the current state $s_{t}\in\mathcal{S}$ as input and outputs an action $a_{t}\in\mathcal{A}$.
-Then the environment will transit to the next state $s_{t+1}\sim\mathcal{T}(\cdot|s_{t},a_{t})$ and the agent will receive a reward $r_{t}=\mathcal{R}(s_{t},a_{t})$.
-The overall objective of RL agent is to learn a policy $\pi:\mathcal{S}\rightarrow A$ that can be used to select actions that maximize the expected cumulative reward over time, i.e.
-$J=\mathbb{E}_{(s_{t},a_{t})\sim\pi}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}].$
-Within the MDP, we optimize a meta-planner policy that interacts with a composite environment consisting of the navigation world $\mathcal{W}$ (filled with obstacles) and a configurable motion planner $f$ parameterized by $\theta$ and receiving sensory inputs $o$.
-The state space, action space, reward function, and termination conditions are defined as follows:
-State space.
-Following [xu2021applr], we assume a local goal $g$ is always reachable and acts as a waypoint on a global path.
-The direction to $g$, $\phi\in[-\pi,\pi]$, relative to the robot’s orientation, provides directional information.
-Thus, the state is $s_{t}=(o_{t},\phi_{t},\theta_{t-1})$, where $o_{t}$ are sensory inputs, $\phi_{t}$ the angle to $g$, and $\theta_{t-1}$ the previous planner configuration.
-Action space.
-The agent’s action $a_{t}=\pi(s_{t})$ updates the planner’s parameters $\theta_{t}$.
-Executing $a_{t}$ applies $\theta_{t}$ to planner $f$, moving the robot and transitioning to the next state $s_{t+1}$.
-Reward.
-The agent receives an immediate reward $r_{t}$ inversely related to its distance to $g$.
-An episode ends when a maximum time $T$ is reached or a termination condition is triggered, with the final reward penalizing timeouts and collisions.
-The goal of the RL agent within the meta-planner is to learn a policy $\pi$ that can select a planner parameter $\theta_{t}$ for $f$ to achieve optimal navigation performance.
-The objective can be formulated as:
-$\max_{\begin{subarray}{c}\pi\end{subarray}}J^{\pi}=\mathbb{E}_{s_{0},\theta_{t}\sim\pi(s_{t}),s_{t+1}\sim\mathcal{T}(s_{t},\theta_{t})}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}r_{t}\Big{]}.$ (1)
-II-D Prioritized Experience Replay
-Prioritized Experience Replay [schaul15prioritized] improves RL by assigning higher probabilities to more informative experiences in the replay buffer, enhancing learning efficiency.
-It selects samples that significantly contribute to the agent’s progress.
-In contrast, our approach focuses on improving data collection quality by reducing sampling bias and promoting better exploration, resulting in a more informative replay buffer.
-Together, these methods complement each other to advance RL efficiency.
-II-E Roadmap Analysis
+Dynamic parameter tuners adjust parameters on-the-fly for different scenarios [xiao2021learning, Binch2020, xiao2020appld, xu2021applr].
+For example, [xiao2020appld] learns a library of parameter sets and switches based on context, while [xu2021applr] uses deep neural network to optimize parameters under various situations.
+II-C Roadmap Analysis
 Existing roadmap analysis [kavraki1998roadmapanalysis, hsu2006foundationroadmap] often relies on static geometric properties to address challenges like narrow passages, using concepts such as visibility and lookout.
@@
 This behavior-guided method allows us to adaptively improve planning and navigation, leading to better performance than static geometric analysis.
-III Sampling Skew of Existing Methods
-APPLR [xu2021applr], the state-of-the-art dynamic meta-planner, exhibits a sampling skew that can lead to inefficiencies and limit performance improvements.
-In this section, we discuss the sampling skew, key insights, and related technical challenges.
+III Training Inefficiency of Existing Methods
+APPLR [xu2021applr], the state-of-the-art dynamic parameter tuner, exhibits a training inefficiency that can lead to limited performance improvements.
+In this section, we discuss the training inefficiency, key insights, and related technical challenges.
 (a) Iteration-1
@@
 (d) Iteration-90
-Figure 3: Sampling skew in existing RL-based Meta-Planner.
+Figure 3: Training inefficiency in existing parameter tuner.
 For convenient visualization, we show 2D projections of the difficult map from Fig.
 2 and omit the robot body.
-The start position is on the left of the map and the goal is on the right.
-Red cross markers are the robot trajectories at each timestamp.
+The yellow arrow denotes the start position and the yellow circle denotes the goal.
+Red cross markers are the robot trajectories at each timestamp (each subfigure contains 3 trajectories to show the distribution).
 On top and right of the maps are the frequency distribution of robot positions $x$ and $y$.
-Meta-Planner Training.
+Parameter Tuner Training.
 Fig.
@@
 2, where the robot navigates obstacles from the start (left) to the goal (right) without collisions and within a time limit.
-The meta-planner adjusts planner’s configurations, i.e.
-actions $a_{t}=\pi(s_{t})$, at a fixed frequency (e.g., 1 Hz) in response to environmental changes, and we record the robot’s positions for reward calculation and policy training.
+The parameter tuner adjusts planner’s configurations, i.e.
+adjustment $a_{t}=\tau(s_{t})$, at a fixed frequency (e.g., 1 Hz) in response to environmental changes, and we record the robot’s positions for reward calculation and tuner training.
 Fig.
-3(a) shows the early-stage challenges with a randomly initialized policy, where the robot often gets stuck (Area A).
+3(a) shows the early-stage challenges with a randomly initialized tuner, where the robot often gets stuck (Area A).
 Over time (Fig.
-3(b)), the policy improves, avoiding early obstacles but still struggling with tight spaces.
+3(b)), the tuner improves, avoiding early obstacles but still struggling with tight spaces.
 As training progresses (Fig.
@@
 3(d)) improves navigation, but the robot still faces difficulties passing through narrow gates (Area D).
-Sampling Skew Phenomenon. From training iterations in Fig. 3, sampling skew can be summarized by 3 observations:
+Training Inefficiency. From training iterations in Fig. 3, training inefficiency can be summarized by 3 observations:
 Observation I: Challenging scenarios require more sampling.
-In difficult situations like dead corners (Fig.
+In situations like dead corners (Fig.
 3(b)), detours (Fig.
-3(c)), and narrow gates (Figures 3(a), 3(d)), the classical planner’s limitations and insufficient sampling of the meta-planner often cause delays, timeouts, or collisions.
+3(c)), and narrow gates (Figures 3(a), 3(d)), the classical planner’s limitations and insufficient sampling of the parameter tuner often cause delays, timeouts, or collisions.
 These critical hard-to-pass areas require significantly more sampling to improve navigation efficiency and safety.
-Observation II: Challenging scenarios have impact on subsequent scenarios.
+Observation II: Challenging scenarios prevent training on subsequent scenarios.
 Hard-to-pass areas make the following scenarios (i.e., hard-to-reach) more difficult to access.
@@
 3(b)) and lower frequency in subsequent regions.
-Observation III: Challenging scenarios tend to be under-fitted.
+Observation III: It is hard and expensive to train on challenging scenarios.
 Certain challenging scenarios remain under-fitted, risking mission success due to two key factors.
 For instance, Area D initially remains hard-to-reach (Fig.
-3(a)-3(c)) and is not effectively trained.
+3(a)-3(c)) and is not accessible to the robot.
 As training progresses (Fig.
-3(d)), Area D becomes hard-to-pass, as the robot reaches it smoothly but takes around 20 steps (about 20 seconds) per trajectory.
-This lengthy process limits efficient training within a fixed budget, making it difficult to fully train Area D.
-Challenges.
-A simple solution to mitigate sampling skew is to focus on training the meta-planner more in under-fitted areas, but there is no automatic way to identify these regions.
-Analyzing static maps alone is insufficient (e.g., detecting narrow parts or sharp turns), as this only considers static information and ignores how the robot dynamically executes and interacts with its environment.
-Our key insight is that hard-to-reach areas stem from preceding hard-to-pass areas.
-If there are no hard-to-pass areas, the robot can easily and safely reach its goal.
-Therefore, our robot behavior-guided method focuses on: (1) increasing sampling in hard-to-pass areas based on Observations I and II, and (2) reducing the time needed to initiate such sampling, aligning with Observation III.
-In Section IV, we describe how we identify these areas to address issues in meta-planner.
-IV Design of Digimon
-We introduce Digimon, consisting of two components: diagnosis (Section IV-B) and mitigation (Section IV-C) of sampling skew phenomenon, as discussed in Section III.
+3(d)), Area D becomes accessible, as the robot can reach it smoothly but takes around 20 steps (about 20 seconds) per trajectory to arrive in Area D.
+This lengthy process is unnecessary but expensive and makes the tuner difficult to sufficiently train on Area D, leading to inefficient training within the given time budget.
+Our Insight.
+Observations I and II highlight the importance of accurately identifying challenging regions and adequately training within them.
+However, Observation III reveals that extensive training in these areas is computationally costly, necessitating efficiency improvements.
+To address these issues, we propose a two-stage approach: (1) leveraging a robot-behavior-guided method to pinpoint critical and challenging regions effectively, and (2) employing targeted up-sampling strategies to reduce the overhead associated with sampling initiation.
+In Section IV, we detail our methodology for identifying these critical regions, directly addressing the inefficiency issue in parameter tuner training.
+IV Design of EffiTune
+We introduce EffiTune (shown in Fig. 4), with two stages: training inefficiency diagnosis (Section IV-B) and mitigation (Section IV-C) for parameter tuners.
 IV-A Overview
 Alg.
-1 illustrates the integration of Digimon within the meta-planner policy training pipeline.
-Components enhanced by Digimon are highlighted in blue, indicating the modifications to the standard meta-planner policy.
-The remaining parts abstract the conventional components of a reinforcement learning training pipeline.
+1 illustrates the integration of EffiTune within the parameter tuner training pipeline.
+Components enhanced by EffiTune are highlighted in blue, indicating the modifications to the standard parameter tuner.
+The remaining parts abstract the conventional components of a deep learning training pipeline.
 To better quantitatively pinpoint the hard-to-pass areas discussed in Section III, we develop Alg.
@@
 Intuitively, high-resistance refers to the observation of abrupt behavioral changes by the robot at these points, suggesting that these areas pose significant uncertainty to the robot’s behavior.
-Algorithm 1 Digimon for Enhancing RL-based Meta-Planner Training
-1:Input: Initial meta-planner policy network $\pi$, high-resistance area $\mathcal{H}$, environment $\mathcal{E}=f\times\mathcal{W}$ for $\pi$ (where $f$ is classical planner and $\mathcal{W}$ is physical map).
-2:Constants: Robot’s initial position ${\tt INIT\_POSE}$, threshold $\lambda$ for sampling initial pose from $\mathcal{H}$, policy training iteration $N$, trajectory collection number $K$, policy update number $L$.
-3:Output: The converged meta-planner policy network $\pi^{*}$.
+Figure 4: Overview of EffiTune Framework
+Algorithm 1 EffiTune for Accelerating Parameter Tuner Training
+1:Input: Initial parameter tuner $\tau$, high-resistance area $\mathcal{H}$, environment $\mathcal{E}=f\times\mathcal{W}$ for $\tau$ (where $f$ is classical planner and $\mathcal{W}$ is physical map).
+2:Constants: Robot’s initial position $\tt INIT\_POSE$}, threshold $\lambda$ for sampling initial pose from $\mathcal{H}$, tuner training iteration NN, trajectory collection number $K$, tuner update number $L$.
+3:Output: The converged parameter tuner $\tau^{*}$.
 4:Set $\mathcal{D}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{P}=\mathcal{H}=\emptyset}$. $\triangleright$ transition buffer $\mathcal{D}$, pose buffer $\mathcal{P}$
-5:for iteration $n=0,...,N$ do
-6:    for $k=0,...,K$ do
+5:for iteration n=0,,Nn=0,...,N do
+6:  for k=0,,Kk=0,...,K do
 7:   Initialize state $s_{0}$ by resetting $\mathcal{E}$.
@@
 9:   Set transition buffer $\mathcal{D}_{k}=\emptyset$, robot pose buffer $\mathcal{P}_{k}=\emptyset$.
-10:        if ${\tt getRandom}($0,1)))) < $\lambda$ and $\mathcal{H}\neq\emptyset$ then
+10:   if ${\tt getRandom}($0,1)) < $\lambda$ and $\mathcal{H}\neq\emptyset$ then
 11:     $p_{0}\leftarrow{\tt getRandom}(\mathcal{H})$
 12:   Set $\tilde{s}=[s_{0}]$ and $\tilde{p}=[p_{0}]$.
-13:        for $t=0,...,T$ do
-14:           Run $\pi$ in $\mathcal{E}$, take action $a_{t}=\pi(s_{t})$.
-15:           Obtain reward $r_{t}$, state $s_{t+1}$ and pose $p_{t+1}$.
+13:   for t=0,,Tt=0,...,T do
+14:     Run $\tau$ in $\mathcal{E}$, take tuning adjustment $a_{t}=\tau(s_{t})$.
+15:     Obtain feedback $r_{t}$, state $s_{t+1}$ and pose $p_{t+1}$.
 16:     $\tilde{s}\leftarrow\tilde{s}.{\tt append}(s_{t+1})$, $\tilde{p}\leftarrow\tilde{p}.{\tt append}(p_{t+1})$.
-17:           if $\pi$ gets ${\tt done}$ signal from $\mathcal{E}$ then,
+17:     if $\tau$ gets $\tt done$} signal from $\mathcal{E}$ then,
 18:      $\mathcal{D}_{k}\leftarrow\mathcal{D}_{k}\cup\{\tilde{s}\}$, $\mathcal{P}_{k}\leftarrow\mathcal{P}_{k}\cup\{\tilde{p}\}$.
@@
 21:  for $\ell=0,...,L$ do
-22:        $\mathcal{B}\leftarrow{\tt getMiniBatch}(\mathcal{D})$triangleright Random mini-batch
-23:        Update policy $\pi$ using $\mathcal{B}$. $\triangleright$ Algorithm-specific
-24:Return the final meta-planner policy $\pi^{*}$.
-IV-B Sampling Skew Diagnosis
+22:   $\mathcal{B}\leftarrow{\tt getMiniBatch}(\mathcal{D})$ $\triangleright$ Random mini-batch
+23:   Update tuner $\tau$ using $\mathcal{B}$. $\triangleright$ Algorithm-specific
+24:Return the final parameter tuner $\tau^{*}$.
+IV-B Training Inefficiency Diagnosis
 High-Resistance Point Identification.
@@
 2, the differences in $x$ and $y$ coordinates between consecutive robot poses are used to calculate trajectory orientations, notated as $\rho_{i}=\arctan(\Delta y_{i}/\Delta x_{i})$ and $\rho_{i+1}=\arctan(\Delta y_{i+1}/\Delta x_{i+1})$.
-Then variation of trajectory orientations $\Delta\rho_{i}=\rho_{i+1}-\rho_{i}$ is then normalized to the range $[-\pi,\pi]$ (in Line 12 of Alg.
+Then variation of trajectory orientations $\Delta\rho_{i}=\rho_{i+1}-\rho_{i}$ is then normalized to the range $\pi,\pi]$ (in Line 12 of Alg.
 2).
@@
 The key intuition behind Alg.
-2 is to capture regions where the meta planner shows high uncertainty in challenging scenarios, indicated by a high variance in the trajectory.
+2 is to capture regions where the parameter tuner shows high uncertainty in challenging scenarios, indicated by a high variance in the trajectory.
 Note that there are two technical aspects that merit particular attention: ❶ Alg.
@@
 This exclusion ensures that any points of high resistance can ultimately reach the goal.
-Allowing the robot to persist in areas that are excessively challenging would cause the planner to become indefinitely stuck, thereby introducing noise into the training of the meta-planner.
+Allowing the robot to persist in areas that are excessively challenging would cause the planner to become indefinitely stuck, thereby introducing noise into the training of the parameter tuner.
 This process allows for the identification of challenging areas where the robot experiences significant difficulties, enabling targeted interventions to improve the planner’s performance in these regions.
+IV-C Training Inefficiency Mitigation
+High-Resistance Area Up-sampling.
+In Lines 10-11 of Alg.
+1, the process of high-resistance area up-sampling is implemented to enhance the robustness and adaptability of the parameter tuner during training.
+This algorithm leverages threshold $\lambda\in[0,1]$ to determine whether to sample the robot’s initial position from the set of high-resistance areas $\mathcal{H}$.
+It ensures that the parameter tuner frequently encounters and learns to navigate through high-resistance regions, thereby improving its ability to handle challenging scenarios.
+By balancing this targeted exploration with a general exploration of the environment, the parameter tuner will train more efficiently and effectively in high-resistance areas.
+Additionally, $\mathcal{H}$ is updated iteratively in Line 20 of Alg.
+1, ensuring that the set of high-resistance areas remains relevant and accurately reflects the current sampling challenges.
+This iterative update is crucial for maintaining the effectiveness of the up-sampling strategy.
+Overall, the up-sampling of high-resistance areas is critical for the parameter tuner’s training as it enhances the planner’s robustness, promotes better generalization to new environments, and addresses training inefficiency.
 Algorithm 2 High-Resistance Area Identification $\tt getHRArea$}
@@
 15:Return the high-resistance area $\mathcal{H}$.
-IV-C Sampling Skew Mitigation
-High-Resistance Area Up-sampling.
-In Lines 10-11 of Alg.
-1, the process of high-resistance area up-sampling is implemented to enhance the robustness and adaptability of the meta-planner during training.
-This algorithm leverages threshold $\lambda\in[0,1]$ to determine whether to sample the robot’s initial position from the set of high-resistance areas $\mathcal{H}$.
-It ensures that the meta-planner frequently encounters and learns to navigate through high-resistance regions, thereby improving its ability to handle challenging scenarios.
-By balancing this targeted exploration with a general exploration of the environment, the meta-planner will train more efficiently and effectively in high-resistance areas.
-Additionally, $\mathcal{H}$ is updated iteratively in Line 20 of Alg.
-1, ensuring that the set of high-resistance areas remains relevant and accurately reflects the current sampling challenges.
-This iterative update is crucial for maintaining the effectiveness of the up-sampling strategy.
-Overall, the up-sampling of high-resistance areas is critical for the meta-planner’s training as it enhances the planner’s robustness, promotes better generalization to new environments, and addresses sampling skew.
 V Evaluation
-Table I: Navigation performance of Digimon and baseline methods under Same-Env setup.
-Metric $\Rightarrow$ Method $\Downarrow$ Easy Medium Difficult
+Table I: Navigation performance of EffiTune and baseline methods under Same-Env setup.
+Metric $\Rightarrow$ Tuner $\Downarrow$ Easy Medium Difficult
 NS $\uparrow$ ATT $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$ NS $\uparrow$ ATT $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$ NS $\uparrow$ ATT $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$
-DWA 20.72 30.01 97.65 1.18 1.18 16.91 33.35 90.59 2.35 7.06 10.61 30.99 53.75 16.25 30.00
-DWA-Fast 25.44 23.60 76.47 23.53 0.00 16.30 30.68 64.71 23.53 11.76 12.43 27.40 50.00 18.75 31.25
+DWA (No Tuner) 20.72 30.01 97.65 1.18 1.18 16.91 33.35 90.59 2.35 7.06 10.61 30.99 53.75 16.25 30.00
+Manual 25.44 23.60 76.47 23.53 0.00 16.30 30.68 64.71 23.53 11.76 12.43 27.40 50.00 18.75 31.25
 BayesOpt 17.88 48.09 74.00 10.00 16.00 15.61 21.84 49.02 21.57 29.41 5.94 58.37 43.75 6.25 50.00
@@
   • NS: Navigation Score. ATT: Actual Traveling Time. SR: Success Rate. CR: Collision Rate. TR: Timeout Rate.
-Table II: Navigation performance of Digimon and baselines under Cross-Env and Cross-Level setups.
+Table II: Navigation performance of EffiTune and baselines under Cross-Env and Cross-Level setups.
 The column names represent the test environment levels: All (across all three levels), Easy, Medium and Difficult.
-Training Env. Metric $\Rightarrow$ Method $\Downarrow$ All Easy Medium Difficult
+Training Env. Metric $\Rightarrow$ Tuner $\Downarrow$ All Easy Medium Difficult
 NS $\uparrow$ ATT $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$ NS $\uparrow$ ATT $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$ NS $\uparrow$ ATT $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$ NS $\uparrow$ ATT $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$
@@
   • NS: Navigation Score. ATT: Actual Traveling Time. SR: Success Rate. CR: Collision Rate. TR: Timeout Rate.
+In evaluation, we answer research questions (RQ) below:
+RQ1: How is EffiTune’s tuning performance compared with baseline tuning methods (under the same time budget)?
+RQ2: Can EffiTune improve parameter tuner’s robustness under different environments and difficulty levels?
+RQ3: How effectiveness is EffiTune’s diagnosis (region identification and failure filtering) for parameter tuners?
+RQ4: Can EffiTune improve the efficiency of training process of parameter tuners?
+RQ5: Can the advantage of EffiTune be illustrated through a case study?
 V-A Experiment Setup
 Simulation & Planners.
-Following existing works [xiao2020appld, xu2021applr], we implement Digimon using the widely-used Gazebo [Gazebo] simulator with a ClearPath Jackal differential drive ground robot as a robot model.
-Due to space limit, we leave the details of the classical planner and the meta-planner that Digimon builds upon for evaluation to Appendix VII.
+Following existing works [xiao2020appld, xu2021applr], we implement EffiTune using the widely-used Gazebo [Gazebo] simulator with a ClearPath Jackal differential drive ground robot as a robot model.
+Due to space limit, we leave the details of the classical planner and the parameter tuner that EffiTune builds upon for evaluation to Appendix I.
 Baselines & Dataset.
-We compare Digimon with two widely used classical planners (DWA [fox1997dwa] and DWA-Fast [dwafast]), one representative static meta-planner, BayesOpt [snoek2012practical] and one SOTA dynamic RL-based meta-planner APPLR [xu2021applr].
+We compare EffiTune with the following tuning setups: (1) classical planners (DWA [fox1997dwa] without parameter tuner; (2) manually-tuned by expert (DWA-Fast [dwafast]); (3) one representative static parameter tuner, BayesOpt [snoek2012practical] and (4) one SOTA dynamic parameter tuner APPLR [xu2021applr].
 We choose the same dataset as APPLR, namely BARN [perille2020benchmarking] dataset, as it includes various maps for robot navigation inside the Gazebo simulator and it has different difficulty levels’ map (i.e., Easy, Medium, and Difficult) to evaluate.
@@
 Comparison Setup.
-We introduce 3 evaluation setups: (1) Same-Env: The meta-planner is trained and tested in the same environment set with predefined but different start positions to avoid overfitting.
-(2) Cross-Env: The meta-planner is trained and tested at the same difficulty level but in different environments.
-(3) Cross-Level: The meta-planner is trained on one difficulty level but tested on a different difficulty level.
-V-B Same-environment Performance
-We evaluate Digimon and four baselines under the Same-Env setup.
+We introduce 3 evaluation setups: (1) Same-Env: The parameter tuner is trained and tested in the same environment set with predefined but different start positions to avoid overfitting.
+(2) Cross-Env: The parameter tuner is trained and tested at the same difficulty level but in different environments.
+(3) Cross-Level: The parameter tuner is trained on one difficulty level but tested on a different difficulty level.
+V-B Same-environment Performance (RQ1)
+We evaluate EffiTune and four baselines under the Same-Env setup.
 DWA, DWA-Fast, and BayesOpt are static planners, so we directly implement their fixed configurations from the training environment and test them on the same environments with random starting positions.
-For APPLR and Digimon, we train the meta-planner policy on the training set and deploy it on the test environments.
+For APPLR and EffiTune, we train the parameter tuner on the training set and deploy it on the test environments.
+Table I shows that EffiTune consistently achieves the highest NS score across all difficulty levels, demonstrating its ability to identify and improve performance in underrepresented scenarios.
+While DWA performs better on SR and CR in easy and medium environments, EffiTune excels in the comprehensive NS metric.
+DWA’s conservative design prioritizes safety over time efficiency, leading to much longer navigation times, as seen in reduced ATT and lower NS scores, making it less practical and ideal for navigation tasks.
+Since NS is a widely-used metric [xu2021applr, xiao2020appld], EffiTune’s higher NS score effectively demonstrates its superiority over both static planners and the SOTA dynamic parameter tuner.
+(a) Trained on Easy Level
+(b) Trained on Medium Level
+(c) Trained on Difficult Level
+Figure 5: Comparison of RandomSampling and EffiTune. Parameter tuners are tested on all 3 levels.
+V-C Cross-environment & Cross-level Performance (RQ2)
+In this section, we compare EffiTune and selected baselines under the Cross-Env and Cross-Level setup, as shown in the diagonal sub-tables in Table II.
+Under the Cross-Env setup, EffiTune consistently achieves the highest NS score across all three environments, indicating its effectiveness in identifying and addressing underrepresented scenarios to enhance planner performance.
+For the Cross-Level experiments, EffiTune excels in generalization by achieving the highest NS scores when the trained parameter tuner is tested across different environmental difficulty levels.
+This demonstrates EffiTune’s robustness and capability to maintain navigation efficacy in new and varying conditions.
+Additionally, when parameter tuners trained in difficult environments are applied to easier ones, EffiTune outperforms even those baselines specifically trained for simpler environments, while this trend is not observed in the baseline methods.
+This result suggests that the parameter tuners trained by our method can learn more sophisticated and adaptable configurations, enabling EffiTune to excel in environments that are less challenging.
+V-D Diagnosis Effectiveness (RQ3)
+To better understand why our diagnosis strategy is effective, we introduce a variant called RandomSampling (RS).
+Unlike EffiTune, which identifies high-resistance areas with Alg.
+2, RS indiscriminately samples from all locations reached by the robot during training without strategic guidance, while all other designs remain consistent with EffiTune.
+Results in Fig.
+5 show that EffiTune significantly outperforms RS, with higher NS scores in all environments.
+The reasons are two-fold.
+First, RS fails to expose the robot to challenging, hard-to-reach scenarios, limiting effective training in these high-resistance areas, which is a crucial aspect for satisfactory navigation performance.
+Second, as shown in Fig.
+2, medium and hard environments have more complicated maps with many hard-to-reach scenarios and obstacles than the easy level.
+Random sampling in such environments often leads to the robot encountering “dead-corner” scenarios – areas that are hard to exit, leading to extended, unrewarding training trajectories.
+Such trajectories will destabilize the training of parameter tuner and degrade navigation performance.
+This analysis verifies the necessity of our high-resistance area up-sampling step in EffiTune.
+By strategically selecting those hard-to-reach areas, we enhance the robot’s ability to perform well across a variety of navigational scenarios.
+We also conduct an ablation study about failure trajectory filtering.
+Table III shows the impact of applying failure filtering on EffiTune’s performance metrics within the same computational budget.
+With filtering enabled, the Navigation Score (NS) improved to 23.23, demonstrating better navigation quality.
+Additionally, the Success Rate (SR) increased significantly from 70.59% to 76.47%, accompanied by a notable reduction in the Collision Rate (CR) from 22.06% to 17.65%.
+This indicates that the filtering process effectively removes noisy or less beneficial training data, enabling the parameter tuner to train more efficiently and enhancing overall robot navigation performance.
+Appendix II shows more results in the comparison between RS and EffiTune.
 Figure 6: Navigation score (NS) of different checkpoints.
-Table III: Performance of Digimon on PPO-based meta-planner.
-PPO + APPLR PPO + Ours
-NS $\uparrow$ 17.58 20.61
-ATT $\downarrow$ 24.69 22.07
-SR $\uparrow$ 55.56 65.28
-CR $\downarrow$ 26.39 15.28
-TR $\downarrow$ 18.06 19.44
-Table IV: Performance of Digimon w/ and wo/ failure filtering.
+V-E Training Efficiency (RQ4)
+In this section, we evaluate the performance of different parameter tuner checkpoints during training under the Cross-Env setup when EffiTune is applied in a medium-level environment.
+Fig.
+6 shows the average NS metric when those checkpoints are evaluated in testing environments.
+EffiTune consistently outperforms APPLR, achieving higher navigation scores at all training stages with the same training steps.
+Furthermore, APPLR requires over four times training steps to achieve comparable performance of EffiTune, and our method ultimately outperforms APPLR by 13.5%.
+These results demonstrate that our method significantly enhances the training efficiency of the parameter tuner and improves the overall navigation performance of the robot.
+Besides NS, we also show more metrics of different parameter tuner checkpoints during training (Appendix III).
+Table III: Ablation Study. Performance of EffiTune with and without applying failure filtering.
 wo/ Filtering w/ Filtering
@@
 TR $\downarrow$ 7.35 5.88
-Table I shows that Digimon consistently achieves the highest NS score across all difficulty levels, demonstrating its ability to identify and improve performance in underrepresented scenarios.
-While DWA performs better on SR and CR in easy and medium environments, Digimon excels in the comprehensive NS metric.
-DWA’s conservative design prioritizes safety over time efficiency, leading to longer navigation times, as seen in reduced ATT and lower NS scores, making it less practical for navigation tasks.
-Since NS is a widely-used metric [xu2021applr, xiao2020appld], Digimon’s higher NS score effectively demonstrates its superiority over both static planners and the SOTA dynamic meta-planner.
-V-C Cross-environment & Cross-level Performance
-In this section, we compare Digimon and selected baselines under the Cross-Env and Cross-Level setup, as shown in the diagonal sub-tables in Table II.
-Under the Cross-Env setup, Digimon consistently achieves the highest NS score across all three environments, indicating its effectiveness in identifying and addressing underrepresented scenarios to enhance planner performance.
-For the Cross-Level experiments, Digimon excels in generalization by achieving the highest NS scores when the trained meta-planner is tested across different environmental difficulty levels.
-This demonstrates Digimon’s robustness and capability to maintain navigation efficacy in new and varying conditions.
-Additionally, when meta-planners trained in difficult environments are applied to easier ones, Digimon outperforms even those baselines specifically trained for simpler environments, while this trend is not observed in the baseline methods.
-This result suggests that the meta-planners trained by our method can learn more sophisticated and adaptable configurations, enabling Digimon to excel in environments that are less challenging.
-V-D Diagnosis Effectiveness
-To better understand why our diagnosis and mitigation strategy is effective, we introduce a variant called RandomSampling (RS).
-Unlike Digimon, which identifies high-resistance areas with Alg.
-2, RS indiscriminately samples from all locations reached by the robot during training without strategic guidance, while all other designs remain consistent with Digimon.
-Results in Fig.
-5 show that Digimon significantly outperforms RS, with higher NS scores in all environments.
-The reasons are two-fold.
-First, RS fails to expose the robot to challenging, hard-to-reach scenarios, limiting effective training in these high-resistance areas, which is a crucial aspect for satisfactory navigation performance.
-Second, as shown in Fig.
-2, medium and hard environments have more complicated maps with many hard-to-reach scenarios and obstacles than the easy level.
-Random sampling in such environments often leads to the robot encountering “dead-corner” scenarios – areas that are hard to exit, leading to extended, unrewarding training trajectories.
-Such trajectories will destabilize the training of the RL policy in the meta-planner and degrade navigation performance.
-This analysis verifies the necessity of our high-resistance area up-sampling step in Digimon.
-By strategically selecting those hard-to-reach areas, we enhance the robot’s ability to perform well across a variety of navigational scenarios.
-Appendix VIII shows more results in the comparison between RS and Digimon.
-(a) Easy
-(b) Medium
-(c) Difficult
-Figure 5: Comparison of RandomSampling and Digimon. The meta-planners are trained on three different difficulty levels and tested across all levels.
-V-E Efficiency
-In this section, we evaluate the performance of different meta-planner policy checkpoints during training under the Cross-Env setup when Digimon is applied in a medium-level environment.
-Fig.
-4 shows the average NS metric when those checkpoints are evaluated in testing environments.
-Digimon consistently outperforms APPLR, achieving higher navigation scores at all training stages with the same training steps.
-Furthermore, APPLR requires over four times training steps to achieve comparable performance of Digimon, and our method ultimately outperforms APPLR by 13.5%.
-These results demonstrate that our method significantly enhances the training efficiency of the RL policy within the meta-planner and improves the overall navigation performance of the robot.
-Besides NS, we also show more metrics of different meta-planner checkpoints during training (in Fig.
-9 in Appendix).
-V-F Plug-and-Play Applicability on Other RL Algorithm
-In Sections V-B and V-C, we have demonstrated that Digimon enhances navigation performance when paired with the off-policy algorithm TD3 [fujimoto2018addressing] based meta-planner in APPLR.
-To further validate the versatility of Digimon, we extended it to another SOTA on-policy RL algorithm, Proximal Policy Optimization (PPO) [ppo].
-For a fair comparison, we replace the TD3 training algorithm in APPLR with PPO.
-Results are shown in Table V-B.
-We can observe that Digimon outperforms baseline methods when applied to PPO, where four out of five metrics get significantly improved, and a negligible drop in timeout rate.
-This indicates that compared to APPLR, our diagnostic method is robust across various training algorithms and not limited to a specific algorithm.
-V-G Other Experiments
-Ablation study. We conduct the ablation study to justify the importance of key components in Digimon and the sensitivity test against the key hyper-parameters. Details can be found in Appendix IX-B.
+V-F Case Study (RQ5)
+In Fig.
+7, we show a comparison before and after applying EffiTune to the parameter tuner.
+By leveraging our diagnosis and up-sampling techniques, the training process no longer always begins from the default position.
+Instead, it starts directly in the critical region, where the robot behavior exhibits signals of struggle and challenge.
+By skipping less beneficial training data and emphasizing the critical samples, we significantly boost the training efficiency of the parameter tuners.
+(a) Original: Iter. 10 (Fig. 3(b))
+(b) With EffiTune: Iter. 10
+(c) Original: Iter. 90 (Fig. 3(d))
+(d) With EffiTune: Iter. 90
+Figure 7: Case Study for Illustration. The yellow arrow denotes the start position and the yellow circle denotes the goal.
+V-G Ablation study
+We conduct ablation study to justify the importance of key components in EffiTune and the sensitivity test against the key hyper-parameters. Details can be found in Appendix IV.
 VI Conclusion
-In this paper, we identify and analyze the challenges of integrating RL with classical planning in meta-planners.
-While this integration enhances adaptability and explainability, it introduces sampling skew due to the conservative nature of classical planners, limiting the RL meta-planners’ exploration and performance in complex environments.
-To address these issues, we propose Digimon, a diagnostic and mitigation framework for RL-based meta-planners.
-Digimon analyzes robot behaviors, identifies exploration bottlenecks and applies a targeted up-sampling strategy to balance the sampling data, enabling more effective learning and exploration.
-Our experiments demonstrate that Digimon improves navigation performance by 13.5%+, enhances robustness in unseen environments, and quadruples training efficiency.
+In this paper, we identify and analyze the training inefficiency of parameter tuners in robot navigation systems.
+To address these issues, we propose EffiTune, a diagnostic and mitigation framework for parameter tuners.
+EffiTune analyzes robot behaviors, identifies exploration bottlenecks and applies a targeted up-sampling strategy to balance the sampling data, enabling more effective learning and exploration.
+Our experiments demonstrate that EffiTune improves navigation performance by 13.5%+, enhances robustness in unseen environments, and quadruples training efficiency.
 Limitations.
-Although Digimon greatly boosts efficiency, further optimization of training duration is possible.
+Although EffiTune greatly boosts efficiency, further optimization of training duration is possible.
 The current pipeline depends on executing the classical planner to collect data, meaning increased complexity of the classical planner can raise training costs.
-Additionally, Digimon is specifically designed for navigation tasks and may not apply to manipulation tasks, which we leave for future work.
+Acknowledgement
+We are grateful to the Center for AI Safety for providing computational resources.
+This work was funded in part by the National Science Foundation (NSF) Awards SHF-1901242, SHF-1910300, Proto-OKN 2333736, IIS-2416835, DARPA VSPELLS - HR001120S0058, ONR N00014-23-1-2081, and Amazon.
+Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
 Appendix
-VII Experiment Setup
+I Experiment Setup
 We run experiments on Ubuntu 20.04, with Intel i9-13900K, 64 GB RAM and Nvidia GPU RTX 2080.
@@
 The onboard Robot Operating System (ROS) $\tt move\_base$} navigation stack (our underlying classical motion planner $f$) uses Dijkstra’s algorithm [dijkstra2022note] to plan a global path and runs DWA [fox1997dwa] as the local classical planner to follow the global path.
-The meta-planner policy $\pi$ is trained to update the planning configurations $\theta$ of the classical planner.
-We maintain the same settings as APPLR, utilizing the DWA classical planner and the same action space $\theta$, including ${\tt max\_vel\_x}$, ${\tt max\_vel\_theta}$, ${\tt vx\_samples}$, ${\tt vtheta\_samples}$, ${\tt occdist\_scale}$, ${\tt pdist\_scale}$, ${\tt gdist\_scale}$, and ${\tt inflation\_radius}$ [^1] https://wiki.ros.org/dwa_local_planner We use the ROS dynamic reconfigure client to dynamically change planner parameters.
+The parameter tuner policy $\pi$ is trained to update the planning configurations $\theta$ of the classical planner.
+We maintain the same settings as APPLR, utilizing the DWA classical planner and the same action space $\theta$, including $\tt max\_vel\_x$}, $\tt max\_vel\_theta$}, $\tt vx\_samples$}, $\tt vtheta\_samples$}, $\tt occdist\_scale$}, $\tt pdist\_scale$}, $\tt gdist\_scale$}, and $\tt inflation\_radius$}.
+[^1] https://wiki.ros.org/dwa_local_planner We use the ROS dynamic reconfigure client to dynamically change planner parameters.
 The global goal $\alpha$ is assigned manually, while $\theta_{0}$ is the default set of parameters provided by the robot manufacturer.
-[^2] http://wiki.ros.org/costmap_2d Table V shows the details of each configuration of the meta-planner.
-Configuration Min Max Meaning
-${\tt max\_vel\_x}$ 0.1 2.0 m/s. Maximum threshold for linear velocity.
-${\tt max\_vel\_theta}$ 0.314 3.14 rad/s. The absolute value of the maximum rotational velocity for the robot.
-${\tt vx\_samples}$ 4 12 The number of samples to use when exploring the x velocity space.
-${\tt vtheta\_samples}$ 8 40 The number of samples to use when exploring the theta velocity space
-${\tt path\_distance\_bias}$ 0.1 0.5 The weight for how much the robot should stay close to the path it was given.
-${\tt goal\_distance\_bias}$ 0.1 2 The weight for how much the robot should attempt to reach its local goal.
-${\tt inflation\_radius}$ 0.1 0.6 Controls how far away the zero cost point is from obstacle.
-Table V: Action Space of Meta-Planner Policy $\pi$.
-VIII Diagnosis Effectiveness
-In this section, we delve deeper into the comparative analysis between RandomSampling and Digimon by applying both methods across environments of varying complexity: Easy, Medium, and Difficult.
+[^2] http://wiki.ros.org/costmap_2d Table IV shows the details of each configuration of the parameter tuner.
+II Diagnosis Effectiveness
+In this section, we delve deeper into the comparative analysis between RandomSampling and EffiTune by applying both methods across environments of varying complexity: Easy, Medium, and Difficult.
 We conducted separate experiments for each difficulty level, which allowed for detailed and direct comparisons.
-The results of these experiments are represented in Figures 67, and 8.
-Consistent with the observations discussed in Section V-D, our Digimon continues to outperform RandomSampling in overall navigation performance.
-This superiority is quantified by a higher Navigation Score (NS), indicating Digimon’s effectiveness in handling complex navigation tasks.
-This comparison not only highlights the strengths of Digimon but also contributes to a better understanding of its operational dynamics across different environmental settings.
+The results of these experiments are represented in Figures 8, 9, and 10.
+Consistent with the observations discussed in Section V-D, our EffiTune continues to outperform RandomSampling in overall navigation performance.
+This superiority is quantified by a higher Navigation Score (NS), indicating EffiTune’s effectiveness in handling complex navigation tasks.
+This comparison not only highlights the strengths of EffiTune but also contributes to a better understanding of its operational dynamics across different environmental settings.
 (a) Train on Easy
@@
 (c) Train on Difficult
-Figure 6: Test on Easy Level
+Figure 8: RandSampling v.s. EffiTune. Test on Easy Level.
 (a) Train on Easy
@@
 (c) Train on Difficult
-Figure 7: Test on Medium Level
+Figure 9: RandSampling v.s. EffiTune. Test on Medium Level.
 (a) Train on Easy
@@
 (c) Train on Difficult
-Figure 8: Test on Difficult Level
-Table VI: Navigation performance under different high-resistance area up-sampling ratios $\lambda$.
+Figure 10: RandSampling v.s. EffiTune. Test on Difficult Level.
+(a) Average Time
+[…]
+(b) Average Reward
+[…]
+(c) Average Success Rate
+Figure 11: Average time, reward, and success rate of EffiTune and APPLR for different checkpoints during training.
+Table V: Navigation performance under different high-resistance area up-sampling ratios $\lambda$.
 Ratio $\lambda$ NS $\uparrow$ ATT $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$
@@
 0.8 11.59 27.04 42.50 17.50 40.00
-Table VII: Navigation performance under different trajectory vector change thresholds $\eta$.
+Table VI: Navigation performance under different trajectory vector change thresholds $\eta$.
 Threshold $\eta$ () NS $\uparrow$ ATT $\downarrow$ SR $\uparrow$ CR $\downarrow$ TR $\downarrow$
@@
 130 17.45 23.70 57.65 14.12 28.24
-(a) Average Time
-[…]
-(b) Average Reward
-[…]
-(c) Average Success Rate
-Figure 9: Average time, reward, and success rate of Digimon and APPLR for different checkpoints during training.
-IX Additional Experiments
-IX-A More Results of Efficiency Comparison
-In Figure 9, we present additional metrics for various meta-planner checkpoints during training, including average traveling time, reward, and success rate.
-It clearly shows that Digimon outperforms APPLR across all three metrics.
-Notably, Digimon requires fewer training steps to achieve shorter traveling times, highlighting the method’s efficiency.
-Similarly, within the same amount of training steps, Digimon can also achieve a higher average reward and success rate.
-All of these three metrics robustly demonstrate that Digimon can achieve superior efficacy compared to the SOTA baseline APPLR.
-IX-B Ablation Study
+III More Results of Efficiency Comparison
+In Figure 11, we present additional metrics for various parameter tuner checkpoints during training, including average traveling time, reward, and success rate.
+It clearly shows that EffiTune outperforms APPLR across all three metrics.
+Notably, EffiTune requires fewer training steps to achieve shorter traveling times, highlighting the method’s efficiency.
+Similarly, within the same amount of training steps, EffiTune can also achieve a higher average reward and success rate.
+All of these three metrics robustly demonstrate that EffiTune can achieve superior efficacy compared to the SOTA baseline APPLR.
+IV Ablation Study
 Sensitivity.
-We conduct ablation studies to evaluate the sensitivity of Digimon against various crucial hyper-parameters: (1) high-resistance area sampling threshold $\lambda$ in Algorithm 1 and (2) high-resistance point identification threshold $\eta$ in Algorithm 2.
+We conduct ablation studies to evaluate the sensitivity of EffiTune against various crucial hyper-parameters: (1) high-resistance area sampling threshold $\lambda$ in Algorithm 1 and (2) high-resistance point identification threshold $\eta$ in Algorithm 2.
 Failure Trajectory Filtering.
-We evaluate the necessity of our failure trajectory filtering in Digimon and its advantages for improving navigation performance.
-The results are shown in Table V-B.
+We evaluate the necessity of our failure trajectory filtering in EffiTune and its advantages for improving navigation performance.
+The results are shown in Table III.
 We observe that without the failure filtering trajectory mechanism, three out of five metrics deteriorate significantly, particularly the collision rate and success rate, which are critical to the robot’s safety.
@@
 1 on the navigation performance.
-Table VIII shows that $\lambda=0.4$ yields the best performance for the navigation system.
+Table II shows that $\lambda=0.4$ yields the best performance for the navigation system.
 Varying $\lambda$ introduces fluctuations across various metrics, while all within a 10% range.
 This demonstrates the necessity of fine-tuning the high-resistance area up-sampling ratio $\lambda$ to balance the proportion of difficult trajectories within the total training set.
-If $\lambda$ is set too high, the training set will contain too many challenging trajectories, complicating the meta-planner’s ability to understand the complete map and complete the task.
-If $\lambda$ is set too low, it will revert to the original training method, hindering the meta-planner’s ability to learn from those under-fitted trajectories.
-In Table VIII, we show the performance metrics of Digimon over different threshold angles.
+If $\lambda$ is set too high, the training set will contain too many challenging trajectories, complicating the parameter tuner’s ability to understand the complete map and complete the task.
+If $\lambda$ is set too low, it will revert to the original training method, hindering the parameter tuner’s ability to learn from those under-fitted trajectories.
+In Table II, we show the performance metrics of EffiTune over different threshold angles.
 While there are some improvements in certain metrics (e.g., NS and CR) at higher threshold angles, the performance data clearly converges around 90° as the optimal threshold for overall system performance.
@@
 Thus, we set $\lambda=0.4$ and $\eta=90$° for all other experiments.
-X Potential Social Impacts
-The development and deployment of Digimon, aimed at enhancing the performance of meta-planner policy in robotic navigation, promises significant positive social benefits.
+V Potential Social Impacts
+The development and deployment of EffiTune, aimed at enhancing the performance of parameter tuner policy in robotic navigation, promises significant positive social benefits.
 By enabling robots to navigate more effectively and reliably, our tool can greatly facilitate the integration of robotics into our daily life.
@@
 Furthermore, in disaster response situations, robots that can navigate challenging terrains could save lives by reaching areas that are inaccessible to humans.
-Overall, by improving the reliability and functionality of meta-planners in robotic navigation systems, Digimon has the potential to make significant contributions to society, improving safety, accessibility, and efficiency in numerous critical areas.
+Overall, by improving the reliability and functionality of parameter tuners in robotic navigation systems, EffiTune has the potential to make significant contributions to society, improving safety, accessibility, and efficiency in numerous critical areas.